# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v2

@@ 7. Toda lattice
$\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\geq 0}\rightarrow\mathrm{U}_{\hbar}(\mathfrak{gl}^{R}_{N})\otimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N^{R}_{+}\xrightarrow{{}^{R}\iota_{N}}N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N^{R}_{+},$
where the first map is (6.4). By naturality of the Miura transform and Theorem˜5.9, the square
@@
by Theorem˜B.3. To check the remaining relation (2.9), we can use ˜7.8 and check it inside
@@
which is a straightforward calculation, for instance, see [BFN19, Appendix B(v)]. Therefore, we have a map
@@
whose composition with the Miura transform (7.3) coincides with the GKLO map. It is surjective by the first part of ˜7.5. By ˜7.8, the target is identified with its image in $^{N}\tilde{\mathcal{A}}_{\hbar}$. Then ˜2.17 gives an isomorphism
@@
In view of Remark˜6.5, the $W$-action on $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/N_{-}\cong{}^{N}\tilde{\mathcal{A}}_{\hbar}$ is
@@
7.1. Comparison with geometry
Recall that the isomorphism $\mathrm{Y}^{0}_{-2N}(\mathfrak{sl}_{2})\cong N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$ is constructed in [BFN19] by identifying the source with the homology of the affine Grassmannian and by using the derived Satake equivalence of [BF08]. In this subsection, we compare the isomorphism of loc. cit. with the one constructed in this paper.
On the topological side, consider the affine Grassmannian $\mathrm{Gr}=\mathrm{GL}_{N}((t))/\mathrm{GL}_{N}[[t]]$. For any dominant coweight $\lambda\in\Lambda^{+}$, denote by $\mathrm{Gr}^{\lambda}=G[[t]]t^{\lambda}G[[t]]$ the corresponding orbit and by $\mathrm{IC}^{\lambda}$ the IC sheaf on the closure $\overline{\mathrm{Gr}^{\lambda}}$. The cohomology $H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\overline{\mathrm{Gr}^{\lambda}},\mathrm{IC}^{\lambda})$ is a $H^{\bullet}_{\mathrm{GL}_{N}\times\mathbf{C}^{*}}(\mathrm{pt})$-bimodule via convolution.
@@ 7.1. Comparison with geometry
Analogously to Section˜5, one can define the negative version of the Kostant-Whittaker reduction
$X\mapsto X/\!\!/_{\psi^{*}}N_{-}:=(X/\mathfrak{n}_{-}^{\psi^{*}})^{N_{-}},\qquad\mathfrak{n}_{-}^{\psi^{*}}=\{\xi-\psi^{*}(\xi)|\xi\in\mathfrak{n}_{-}\},$
for any Harish-Chandra bimodule $X$. Recall that the coefficients of the quantum minor $T^{1\ldots N}_{1\ldots N}(u)$ from (1.11) generate the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$. Let us identify it with the generating function of the equivariant Chern classes of the dual representation, see Remark˜7.15:
(7.4) $\mathbf{Z}_{\mathfrak{gl}_{N}}\cong H^{\bullet}_{G}(\mathrm{pt}),\qquad T^{1\ldots N}_{1\ldots N}(u)\mapsto\sum_{i=0}^{N}c_{i}^{\mathrm{GL}_{N}}[(\mathbf{C}^{N})^{*}]u^{N-i}.$
For every dominant weight $\lambda\in\Lambda$, the derived Satake equivalence of [BF08] gives an isomorphism of $\mathbf{Z}_{\mathfrak{gl}_{N}}$-bimodules
(7.5) $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes V_{\lambda}/\!\!/_{\psi^{*}}N_{-}\xrightarrow{\sim}H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\overline{\mathrm{Gr}^{\lambda}},\mathrm{IC}^{\lambda}),$
@@
(7.6) $N^{L}_{-}\,\backslash\!\!\backslash_{-\psi^{*}}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi^{*}}N_{-}^{R}\xrightarrow{\sim}H^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}_{\bullet}(\mathrm{Gr}),$
@@
${\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes V_{\lambda}/\!\!/_{\psi^{*}}\,N^{R}_{-}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\overline{\mathrm{Gr}^{\lambda}},\mathrm{IC}^{\lambda})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N^{L}_{-}\,\backslash\!\!\backslash_{-\psi^{*}}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi^{*}}N_{-}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}_{\bullet}(\mathrm{Gr})}$
@@
$\mathrm{GL}_{N}\rightarrow\mathrm{GL}_{N},\qquad g\mapsto(g^{-1})^{T}.$
@@
(7.7) $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}\xrightarrow{\sim}N^{L}_{-}\,\backslash\!\!\backslash_{-\psi^{*}}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi^{*}}N_{-}^{R}.$
In particular, by ˜1.2 and (7.4), we have
(7.8) $\mathbf{Z}_{\mathfrak{gl}_{N}}[u]\ni A(u)\mapsto\sum_{i=0}^{N}c_{i}^{\mathrm{GL}_{N}}[(\mathbf{C}^{N})^{*}]u^{N-i},$
where $A(u)$ is the series (1.11). For brevity, we will thus denote $A_{i}=c_{i}^{\mathrm{GL}_{N}}[(\mathbf{C}^{N})^{*}]$.
Recall description of $\mathrm{Gr}^{\lambda}$ in terms of lattices, i.e. free $\mathbf{C}[[t]]$-modules of rank $N$ inside $\mathbf{C}((t))^{N}$ [Zhu17]. We will only consider a simple case $\lambda=\omega_{1}^{*}$ (and $\lambda=\omega_{1}$ in Appendix˜B), where $\omega_{1}^{*}=(0,\ldots,0,-1)$. Then
$\mathrm{Gr}^{\omega_{1}^{*}}=\overline{\mathrm{Gr}^{\omega_{1}^{*}}}=\{L_{0}\subset L_{1}\subset z^{-1}L_{0}|\dim L_{1}/L_{0}=1\}\cong\mathbb{P}^{N-1}=\mathbb{P}(\mathbf{C}^{N}),$
@@
(7.9) $\tilde{\varpi}_{N}\colon\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\ltimes S^{\bullet}\mathbf{C}^{N}\xrightarrow{\sim}\mathrm{U}_{\hbar}(\mathfrak{gl}^{R}_{N})\ltimes S^{\bullet}(\mathbf{C}^{N})^{*},\qquad E_{ij}^{R}\mapsto-E_{ji}^{R},\qquad v_{i}\mapsto\phi_{i},$
@@
(7.10) ${\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{\varpi}_{N}}$${N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N^{L}_{-}\,\backslash\!\!\backslash_{-\psi^{*}}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi^{*}}N_{-}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes(\mathbf{C}^{N})^{*}/\!\!/_{\psi^{*}}\,N_{-}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\sim}$${H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr}^{\omega_{1}^{*}},\underline{\mathbf{C}}_{\mathrm{Gr}^{\omega_{1}^{*}}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}_{\bullet}(\mathrm{Gr})}$
@@
(7.11) $\mathcal{V}_{i}\rightarrow\mathrm{Gr}^{\omega_{1}^{*}},\qquad\mathcal{V}_{i}|_{L_{0}\subset L_{1}}=z^{-1}L_{i}/L_{i}$
@@
(7.12) $A^{(i)}(u)=\sum_{j=0}^{N}u^{N-j}A^{(i)}_{j},\qquad A^{(i)}_{j}=(-1)^{j}c_{j}^{\mathrm{GL}_{N}}(\mathcal{V}_{i-1}),$
for $i=1,2$. For instance, under identification $\mathrm{Gr}^{\omega_{1}^{*}}\cong\mathbb{P}^{N-1}$, we have $A^{(1)}(u)=A(u)$ from (7.8). The left (resp. right) action of $H^{\bullet}_{G}(\mathrm{pt})$ on $H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr}^{\omega_{1}^{*}})$ is via the coefficients of $A^{(1)}(u)$ (resp. $A^{(2)}(u)$). Also, consider the line bundle
@@
$H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr}^{\omega_{1}^{*}},\underline{\mathbf{C}}_{\mathrm{Gr}^{\omega_{1}^{*}}})=H^{\bullet}_{\mathrm{GL}_{N}\times\mathbf{C}^{*}}(\mathbb{P}^{N-1})\cong\mathbf{C}[\hbar][A_{1},\ldots,A_{N}][c^{\mathrm{GL}_{N}}_{1}(\mathcal{S})]/(A(c^{\mathrm{GL}_{N}}_{1}(\mathcal{S}))).$
@@
(7.13) $c_{1}^{G}(\mathcal{S})=c_{1}^{\mathrm{GL}_{N}}(\mathcal{S})+\hbar.$
@@
By [BF08, Lemma 4], the Kostant-Whittaker reduction functor is a full embedding when restricted to free Harish-Chandra bimodules. Therefore, we can lift $\phi$ to an automorphism of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}$ which induces an automorphism of the parabolic restriction $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\otimes\mathbf{C}^{N}/\!\!/N_{-}$ as of $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-bimodules. According to the decomposition (4.2), it is isomorphic to a direct sum of $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-bimodules
@@
where $\bar{v}_{i}$ are vectors (4.10). Since $\phi$ was defined on the non-generic bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}$, the automorphism acts as $\bar{v}_{i}\mapsto f_{i}(w)\bar{v}_{i}$ for some polynomial function $f_{i}(w)\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N})$ with a polynomial inverse, therefore, it should be constant $f_{i}(w)=c_{i}$. Moreover, since $\phi$ commutes with the diagonal action of the Weyl group $W=S_{N}\subset\mathrm{GL}_{N}$, we have $c_{i}=c_{j}$ for any $i,j$. Hence $\phi=c\cdot\mathrm{Id}$ by ˜4.5. ∎
@@
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N}^{R})\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}^{R}\cong\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega]/(A(\Omega))\xrightarrow{\sim}\mathbf{C}[\hbar][A_{1},\ldots,A_{N}][c^{G}_{1}(\mathcal{S})]/(A(c^{G}_{1}(\mathcal{S})))$
composed with Proposition˜5.5 is given by $\Omega^{k}\mapsto(-1)^{N-1}c^{\mathrm{GL}_{N}}_{1}(\mathcal{S})^{k}$ for all $k$.
@@
The left $\mathbf{Z}_{\mathfrak{gl}_{N}}$-module structures are compatible by (7.8). Let us show that this map intertwines the right actions. Following the proof of [CK18, Lemma 8.13], consider two short exact sequences on $\mathrm{Gr}^{\omega_{1}^{*}}$:
@@
$c^{G}(\mathcal{V}_{0})=(u+c^{G}_{1}(\mathcal{S}))c^{G}(z^{-1}L_{0}/L_{1})=(u+c^{G}_{1}(\mathcal{S}))\frac{c^{G}(\mathcal{V}_{1})}{c^{G}(z^{-1}L_{1}/z^{-1}L_{0})}=(u+c^{G}_{1}(\mathcal{S}))\frac{c^{G}(\mathcal{V}_{1})}{u+c^{G}_{1}(\mathcal{S})-\hbar},$
@@
which agrees with the formula of Proposition˜4.7.
Therefore, this map is indeed an isomorphism of $\mathbf{Z}_{\mathfrak{gl}_{N}}$-bimodules such that $v_{1}\mapsto(-1)^{N-1}\in H^{\bullet}_{G}(\mathbb{P}^{N-1})$. The equivalence of Bezrukavnikov-Finkelberg gives an a priori different isomorphism with the same normalization condition by [Fin+18, Lemma 2.13]. However, it should coincide with the former by ˜7.12. ∎
Finally, recall the truncated shifted Yangian of ˜2.17. By [BFN19, Theorem B.18], there is a surjective homomorphism
@@
(7.14) $\begin{split}d_{1}(u)&\mapsto\sum_{p=0}^{N}u^{N-p}c_{p}^{\mathrm{GL}_{N}}[(\mathbf{C}^{N})^{*}],\\ x_{-}(u)&\mapsto\frac{1}{u-c^{G}_{1}(\mathcal{Q})-\hbar}\cap[\mathrm{Gr}^{\omega_{1}}],\\ x_{+}(u)&\mapsto(-1)^{N}\frac{1}{u-c^{G}_{1}(\mathcal{S})-\hbar}\cap[\mathrm{Gr}^{\omega_{1}^{*}}],\end{split}$
Here $\mathrm{Gr}^{\omega_{1}}]\in H_{\bullet}^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr})$ is the fundamental class of the embedding $\mathrm{Gr}^{\omega_{1}}\hookrightarrow\mathrm{Gr}$, similarly for $\mathrm{Gr}^{\omega_{1}^{*}}]$.
@@
is commutative. Here, the left vertical arrow is Theorem˜7.9 and the lower horizontal one is the composition of (7.7) and the derived Satake equivalence (7.6).

### v4

@@ 8 Toda lattice
$\mathrm{Y}_{-N,N}(\mathfrak{gl}_{2})^{\geq 0}\rightarrow\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}^{R}_{N}\bigr)\otimes S^{\bullet}\mathbf{C}^{N}/\!\!/_{\psi}\,N^{R}_{+}\xrightarrow{{}^{R}\iota_{N}}N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N^{R}_{+},$
where the first map is (7.4). By naturality of the Miura transform and Theorem 6.9, the square
@@
by Theorem B.3. To check the remaining relation (3.9), we can use Lemma 8.9 and check it inside
@@
which is a straightforward calculation, for instance, see [4, Appendix B (v)]. Therefore, we have a map
@@
whose composition with the Miura transform (8.4) coincides with the GKLO map. It is surjective by the first part of Lemma 8.6. By Lemma 8.9, the target is identified with its image in $^{N}\!\tilde{\mathcal{A}}_{\hbar}$. Then Definition 3.17 gives an isomorphism
@@
In view of Remark 7.5, the $W$-action on $N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/N_{-}\cong{}^{N}\!\tilde{\mathcal{A}}_{\hbar}$ is
@@
8.1 Comparison with geometry
Recall that the isomorphism $\mathrm{Y}^{0}_{-2N}(\mathfrak{sl}_{2})\cong N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}$ is constructed in [4] by identifying the source with the homology of the affine Grassmannian and by using the derived Satake equivalence of [3]. In this subsection, we compare the isomorphism of loc. cit. with the one constructed in this paper.
On the topological side, consider the affine Grassmannian $\mathrm{Gr}=\mathrm{GL}_{N}((t))/\mathrm{GL}_{N}[[t]]$. For any dominant coweight $\lambda\in\Lambda^{+}$, denote by $\mathrm{Gr}^{\lambda}=G[[t]]t^{\lambda}G[[t]]$ the corresponding orbit and by $\mathrm{IC}^{\lambda}$ the IC sheaf on the closure $\overline{\mathrm{Gr}^{\lambda}}$. The cohomology $H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}\bigl(\overline{\mathrm{Gr}^{\lambda}},\mathrm{IC}^{\lambda}\bigr)$ is a $H^{\bullet}_{\mathrm{GL}_{N}\times\mathbf{C}^{*}}(\mathrm{pt})$-bimodule via convolution.
@@ 8.1 Comparison with geometry
Analogously to Section 6, one can define the negative version of the Kostant–Whittaker reduction
$X\mapsto X/\!\!/_{\psi^{*}}N_{-}:=\bigl(X/\mathfrak{n}_{-}^{\psi^{*}}\bigr)^{N_{-}},\qquad\mathfrak{n}_{-}^{\psi^{*}}=\{\xi-\psi^{*}(\xi)\mid\xi\in\mathfrak{n}_{-}\},$
for any Harish-Chandra bimodule $X$. Recall that the coefficients of the quantum minor $T^{1\dots N}_{1\dots N}(u)$ from (2.8) generate the center $\mathbf{Z}_{\mathfrak{gl}_{N}}$. Let us identify it with the generating function of the equivariant Chern classes of the dual representation, see Remark 8.16,
$\mathbf{Z}_{\mathfrak{gl}_{N}}\cong H^{\bullet}_{G}(\mathrm{pt}),\qquad T^{1\dots N}_{1\dots N}(u)\mapsto\sum_{i=0}^{N}c_{i}^{\mathrm{GL}_{N}}\bigl[\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr]u^{N-i}.$ (8.5)
For every dominant weight $\lambda\in\Lambda$, the derived Satake equivalence of [3] gives an isomorphism of $\mathbf{Z}_{\mathfrak{gl}_{N}}$-bimodules
$\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes V_{\lambda}/\!\!/_{\psi^{*}}N_{-}\xrightarrow{\sim}H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}\bigl(\overline{\mathrm{Gr}^{\lambda}},\mathrm{IC}^{\lambda}\bigr),$
@@
$N^{L}_{-}\,\backslash\!\!\backslash_{-\psi^{*}}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi^{*}}N_{-}^{R}\xrightarrow{\sim}H^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}_{\bullet}(\mathrm{Gr}),$ (8.6)
@@
${\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}_{N}^{R}\bigr)\otimes V_{\lambda}/\!\!/_{\psi^{*}}\,N^{R}_{-}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}\bigl(\overline{\mathrm{Gr}^{\lambda}},\mathrm{IC}^{\lambda}\bigr)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N^{L}_{-}\,\backslash\!\!\backslash_{-\psi^{*}}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi^{*}}N_{-}^{R}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}_{\bullet}(\mathrm{Gr})}$
@@
$\mathrm{GL}_{N}\rightarrow\mathrm{GL}_{N},\qquad g\mapsto\bigl(g^{-1}\bigr)^{\mathsf{T}}.$
@@
$N^{L}_{+}\,\backslash\!\!\backslash_{-\psi}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi}\,N_{+}^{R}\xrightarrow{\sim}N^{L}_{-}\,\backslash\!\!\backslash_{-\psi^{*}}\,\mathrm{D}_{\hbar}(\mathrm{GL}_{N})/\!\!/_{\psi^{*}}N_{-}^{R}.$ (8.7)
In particular, by Lemma 2.2 and (8.5), we have
$\mathbf{Z}_{\mathfrak{gl}_{N}}[u]\ni A(u)\mapsto\sum_{i=0}^{N}c_{i}^{\mathrm{GL}_{N}}\bigl[\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr]u^{N-i},$ (8.8)
where $A(u)$ is the series (2.8). For brevity, we will thus denote $A_{i}=c_{i}^{\mathrm{GL}_{N}}\bigl[\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr]$. Recall description of $\mathrm{Gr}^{\lambda}$ in terms of lattices, i.e., free $\mathbf{C}[[t]]$-modules of rank $N$ inside $\mathbf{C}((t))^{N}$ [50]. We will only consider a simple case $\lambda=\omega_{1}^{*}$ (and $\lambda=\omega_{1}$ in Appendix B), where $\omega_{1}^{*}=(0,\dots,0,-1)$. Then
$\mathrm{Gr}^{\omega_{1}^{*}}=\overline{\mathrm{Gr}^{\omega_{1}^{*}}}=\bigl\{L_{0}\subset L_{1}\subset z^{-1}L_{0}\mid\dim L_{1}/L_{0}=1\bigr\}\cong\mathbb{P}^{N-1}=\mathbb{P}\bigl(\mathbf{C}^{N}\bigr),$
@@
$\tilde{\varpi}_{N}\colon\ \mathrm{U}_{\hbar}\bigl(\mathfrak{gl}_{N}^{R}\bigr)\ltimes S^{\bullet}\mathbf{C}^{N}\xrightarrow{\sim}\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}^{R}_{N}\bigr)\ltimes S^{\bullet}\bigl(\mathbf{C}^{N}\bigr)^{*},\qquad E_{ij}^{R}\mapsto-E_{ji}^{R},\qquad v_{i}\mapsto\phi_{i},$ (8.9)
@@
[…]
@@
$\mathcal{V}_{i}\rightarrow\mathrm{Gr}^{\omega_{1}^{*}},\qquad\mathcal{V}_{i}|_{L_{0}\subset L_{1}}=z^{-1}L_{i}/L_{i}$ (8.11)
@@
$A^{(i)}(u)=\sum_{j=0}^{N}u^{N-j}A^{(i)}_{j},\qquad A^{(i)}_{j}=(-1)^{j}c_{j}^{\mathrm{GL}_{N}}(\mathcal{V}_{i-1}),$ (8.12)
for $i=1,2$. For instance, under identification $\mathrm{Gr}^{\omega_{1}^{*}}\cong\mathbb{P}^{N-1}$, we have $A^{(1)}(u)=A(u)$ from (8.8). The left (resp. right) action of $H^{\bullet}_{G}(\mathrm{pt})$ on $H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}\bigl(\mathrm{Gr}^{\omega_{1}^{*}}\bigr)$ is via the coefficients of $A^{(1)}(u)$ (resp. $A^{(2)}(u)$). Also, consider the line bundle
@@
$H^{\bullet}_{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}\bigl(\mathrm{Gr}^{\omega_{1}^{*}},\underline{\mathbf{C}}_{\mathrm{Gr}^{\omega_{1}^{*}}}\bigr)$ ${}=H^{\bullet}_{\mathrm{GL}_{N}\times\mathbf{C}^{*}}\bigl(\mathbb{P}^{N-1}\bigr)$
${}\cong\mathbf{C}[\hbar][A_{1},\dots,A_{N}]\bigl[c^{\mathrm{GL}_{N}}_{1}(\mathcal{S})\bigr]/\bigl(A\bigl(c^{\mathrm{GL}_{N}}_{1}(\mathcal{S})\bigr)\bigr).$
@@
$c_{1}^{G}(\mathcal{S})=c_{1}^{\mathrm{GL}_{N}}(\mathcal{S})+\hbar.$ (8.13)
@@
By [3, Lemma 4], the Kostant–Whittaker reduction functor is a full embedding when restricted to free Harish-Chandra bimodules. Therefore, we can lift $\phi$ to an automorphism of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}$ which induces an automorphism of the parabolic restriction $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})^{\mathrm{gen}}\otimes\mathbf{C}^{N}/\!\!/N_{-}$ as of $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-bimodules. According to the decomposition (5.1), it is isomorphic to a direct sum of $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$-bimodules
@@
where $\bar{v}_{i}$ are vectors (5.9). Since $\phi$ was defined on the non-generic bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes\mathbf{C}^{N}$, the automorphism acts as $\bar{v}_{i}\mapsto f_{i}(w)\bar{v}_{i}$ for some polynomial function $f_{i}(w)\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N})$ with a polynomial inverse, therefore, it should be constant $f_{i}(w)=c_{i}$. Moreover, since $\phi$ commutes with the diagonal action of the Weyl group $W=S_{N}\subset\mathrm{GL}_{N}$, we have $c_{i}=c_{j}$ for any $i$, $j$. Hence $\phi=c\cdot\mathrm{Id}$ by Lemma 5.5. ∎
@@
$\mathrm{U}_{\hbar}\bigl(\mathfrak{gl}_{N}^{R}\bigr)\otimes\mathbf{C}^{N}/\!\!/_{\psi}\,N_{+}^{R}\cong\mathbf{Z}_{\mathfrak{gl}_{N}}[\Omega]/(A(\Omega))\xrightarrow{\sim}\mathbf{C}[\hbar][A_{1},\dots,A_{N}]\bigl[c^{G}_{1}(\mathcal{S})\bigr]/\bigl(A\bigl(c^{G}_{1}(\mathcal{S})\bigr)\bigr)$
composed with Proposition 6.5 is given by $\Omega^{k}\mapsto(-1)^{N-1}c^{\mathrm{GL}_{N}}_{1}(\mathcal{S})^{k}$ for all $k$.
@@
The left $\mathbf{Z}_{\mathfrak{gl}_{N}}$-module structures are compatible by (8.8). Let us show that this map intertwines the right actions. Following the proof of [5, Lemma 8.13], consider two short exact sequences on $\mathrm{Gr}^{\omega_{1}^{*}}$:
@@
$c^{G}(\mathcal{V}_{0})$ ${}=\bigl(u+c^{G}_{1}(\mathcal{S})\bigr)c^{G}\bigl(z^{-1}L_{0}/L_{1}\bigr)=\bigl(u+c^{G}_{1}(\mathcal{S})\bigr)\frac{c^{G}(\mathcal{V}_{1})}{c^{G}\bigl(z^{-1}L_{1}/z^{-1}L_{0}\bigr)}$
${}=\bigl(u+c^{G}_{1}(\mathcal{S})\bigr)\frac{c^{G}(\mathcal{V}_{1})}{u+c^{G}_{1}(\mathcal{S})-\hbar},$
@@
which agrees with the formula of Proposition 5.7. Therefore, this map is an isomorphism of $\mathbf{Z}_{\mathfrak{gl}_{N}}$-bimodules such that $v_{1}\mapsto(-1)^{N-1}\in H^{\bullet}_{G}\bigl(\mathbb{P}^{N-1}\bigr)$. The equivalence of Bezrukavnikov–Finkelberg gives an a priori different isomorphism with the same normalization condition by [14, Lemma 2.13]. However, it should coincide with the former by Lemma 8.13. ∎
Finally, recall the truncated shifted Yangian of Definition 3.17. By [4, Theorem B.18], there is a surjective homomorphism
@@
$d_{1}(u)\mapsto\sum_{p=0}^{N}u^{N-p}c_{p}^{\mathrm{GL}_{N}}\bigl[\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr],$
$x_{-}(u)\mapsto\frac{1}{u-c^{G}_{1}(\mathcal{Q})-\hbar}\cap[\mathrm{Gr}^{\omega_{1}}],$
$x_{+}(u)\mapsto(-1)^{N}\frac{1}{u-c^{G}_{1}(\mathcal{S})-\hbar}\cap\bigl[\mathrm{Gr}^{\omega_{1}^{*}}\bigr],$ (8.14)
Here $\mathrm{Gr}^{\omega_{1}}]\in H_{\bullet}^{\mathrm{GL}_{N}[[t]]\rtimes\mathbf{C}^{*}}(\mathrm{Gr})$ is the fundamental class of the embedding $\mathrm{Gr}^{\omega_{1}}\hookrightarrow\mathrm{Gr}$, similarly for $\bigl[\mathrm{Gr}^{\omega_{1}^{*}}\bigr]$.
@@
is commutative. Here, the left vertical arrow is Theorem 8.10 and the lower horizontal one is the composition of (8.7) and the derived Satake equivalence (8.6).